Jose and Maria work at a restaurant. Jose can make either 10 pancakes or 4 waffies: Marila can make either 8 pancakes or 2 waffes,According to this A. 4 wames. B. 2.5 wames. О С. 1Awame. O D. 2/5 watfle.
Answer #1

Opportunity cost of making one pancake for Jose = No. of waffles that can be produced by Jose/ No. of pancakes that can be produced by Jose

Opportunity cost = 4/10 = 2/5 waffle (option D)